Andile deposits R900 into a savings account paying 6,5% interest per year, compounded quarterly. After three and a half years he withdraws R1 000 from the account and deposits it into a second account paying 11% simple interest per year.


A number of years after Andile deposited the R1 000 into the second account, the accrued amount in the second account is R1 605. The time (correct to two decimal places) the money was invested for, is


[1] 4,36 years.

[2] 5,50 years.

[3] 9,31 years.

[4] 4,53 years.

[5] none of the above


1
Expert's answer
2021-05-24T19:02:56-0400

I = PRT/100

I = amount - principal

I= R1605 - R1000

=605.

Therefore, 605 = PRT/100

(605*100) = (1000*11)T

T = 60500/11000

= 5.50 years